OpenStudy (anonymous):

2^x=18 lo10^6x=1 solve 5^2x=125

OpenStudy (anonymous):

do you want all of these solved or?

OpenStudy (anonymous):

yes plz

OpenStudy (anonymous):

we are solving for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for the last one all you really have to do is divide to get x to equal 5

OpenStudy (anonymous):

im working on the others

OpenStudy (anonymous):

can you rewrite the second one, im a little confused on what it actually says

OpenStudy (anonymous):

log4 2=x

OpenStudy (anonymous):

x=1/2

OpenStudy (anonymous):

so number 2 is 1/2

OpenStudy (anonymous):

log4 2=x is what equals 1/2

OpenStudy (anonymous):

kk ty on that one i am really bad with math

OpenStudy (anonymous):

FLVS?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

If you ever need help, there's this site that will help you get the answer right away Wolframalpha.com you'd have to type your log questions like this log_4 2=x

OpenStudy (anonymous):

it will show you the graphs, solutions, the steps, simplified versions, remainders, and possible zeros to all your questions

OpenStudy (anonymous):

wow ty alot he helped me with number one and it should help me wiv the other u are amazingly sweet for giving me that website ty ty ty ty

OpenStudy (anonymous):

first one was 2^x=1/8 lol my mistake lol